A clothing store sells men's shirts at $152 and $205 per shirt the store sells 32 shirts and tickets in $5659 how many shirts at each price does the store sells?
Answer by richwmiller(17219)

x+y=32,
152*x+205*y=5659
x=32-y
152*(32-y)+205*y=5659
4864-152y+205*y=5659
53*y=795
y=15
x=32-y
x=17 at 152 and y=15 at 205
check
17+15=32
152*x+205*y=5659
152*17+205*15=5659
2584+3075=5659
5659=5659
ok
codetick